OpenStudy (anonymous):

Evaluate: limit as x--> infty of (sqrt{x^2+9x+1}-x)

OpenStudy (anonymous):

\[\sqrt{x^2+9x+1} -x\] ?

OpenStudy (anonymous):

is that what you mean

OpenStudy (anonymous):

\[\frac{ (\sqrt{x^2+9x+1} -x) (\sqrt{x^2+9x+1} +x)}{\sqrt{x^2+9x+1} +x}\]

OpenStudy (anonymous):

\[= \frac{ (x^2+9x+1)-(x)^2}{\sqrt{ x^2(1 +\frac{9}{x} + \frac{1}{x^2} )} +x}\]

OpenStudy (anonymous):

Ok, so just multiply by the conjugate then combine?

OpenStudy (anonymous):

pretty much. It's a standard procedure for anything of that form.

OpenStudy (anonymous):

\[= \frac{9x+1}{x [ \sqrt{1+\frac{9}{x}+\frac{1}{x^2}} +1] }\]

OpenStudy (anonymous):

\[= \frac{ x (9 + \frac{1}{x} ) }{x[\sqrt{1+\frac{9}{x}+\frac{1}{x^2} } +1} = \frac{9}{2}\]

OpenStudy (anonymous):

That last part should say "limit ____" = (9/2) , but I couldn't be bothered writting it.

OpenStudy (anonymous):

Just incase you don't get the second last step. Remember lim x-> infinity (1/x^n) , where n>=1 equals zero

OpenStudy (anonymous):

so you get \[\frac{9 + (0)}{ \sqrt{1+(0)+(0)} +1} = \frac{9}{2}\]